OpenStudy (anonymous):

What is the equation of the circle which is tangent to the line 3x-4y=32, and the center is at point (0,7)

OpenStudy (anonymous):

@ganeshie8 pls help sir

ganeshie8 (ganeshie8):

equation of circle whose center is \((h,k)\) and radius is \(r\) : \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

@ganeshie8 what will happen next?

ganeshie8 (ganeshie8):

familiar with `point to line` distance formula ?

ganeshie8 (ganeshie8):

you may use that formula to find the `radius`, which is the distance between `center` and the `tangent line`

ganeshie8 (ganeshie8):

line : 3x-4y-32 = 0 point : (0, 7) radius = \(\large \dfrac{|3(0) - 4(7) - 32|}{\sqrt{3^2+4^2}} = ?\)

OpenStudy (anonymous):

\[\frac{ 60 }{ \sqrt{25} }\]

ganeshie8 (ganeshie8):

\[\large \dfrac{|3(0) - 4(7) - 32|}{\sqrt{3^2+4^2}} = \dfrac{60}{\sqrt{25}} = \dfrac{60}{5} = 12\]

OpenStudy (anonymous):

sorry divided the wrong number. xD

ganeshie8 (ganeshie8):

:) plugin center and radius in the equation of circle @Faye281 still here ?

OpenStudy (anonymous):

@ganeshie8 would (x-0)^2+(y-7)^2=144

OpenStudy (anonymous):

x^2+y^2-14y+49=144?

OpenStudy (anonymous):

@ganeshie8 sorry. Afk

ganeshie8 (ganeshie8):

that looks good ^

OpenStudy (anonymous):

@ganeshie8 sorry :( x^2+y^2-14y+49=144? is the answer? aren't we going to transfer 144 to the left equation?